The measure of a supplement of an angle is four more than three times the angle. How much is the angle and the supplement. I tried setting it up like
4+3x=180 then I am lost. Please help.

Answer by Theo(2967)

supplement of an angle = 180 minus the angle.

your formula is:

(180 - x) = 3*x + 4 assuming x is the angle.

solve for x to get:

(180 - x) = 3*x + 4 becomes:
180 - x = 3*x + 4
add x to both sides of the equation and subtract 4 from both sides of the eqution to getr:
180-4 = 3x+x
combine like terms to get:
176 = 4x
divide both sides of the equation by 4 to get:
44 = x
this makes the supplement of the angle = 180-x = 180-44 = 136

supplement of the angle is equal to 3 times the angle + 4.

3*44 = 132 + 4 = 136 which is the supplement of the angle so the formula is good and the answer is:

angle = 44 degrees.

supplement of the angle is 136 degrees




problems states that supplement of the angle = 3 * the angle